The stockholders’ equity section of the balance sheet for Mann Equipment Co. at December 31, Year 1, is as follows:

Stockholders’ Equity
Paid-in capital
Preferred stock, ? par value, 6% cumulative,
100,000 shares authorized, 10,000 shares issued and outstanding
$ 200,000
Common stock, $10 stated value, 200,000 shares
authorized, 100,000 shares issued and outstanding
1,000,000
Paid-in capital in excess of par—Preferred 25,000
Paid-in capital in excess of stated value—Common 500,000
Total paid-in capital 1,725,000
Retained earnings 420,000
Total stockholders’ equity $ 2,145,000


Note: The market value per share of the common stock is $42, and the market value per share of the preferred stock is $26.

Required
a. What is the par value per share of the preferred stock?
b. What is the dividend per share on the preferred stock? (Round your answer to 2 decimal places.)
c. What was the average issue price per share (price for which the stock was issued) of the common stock?
e. If Mann declares a 2-for-1 stock split on the common stock, how many shares will be outstanding after the split? What amount will be transferred from the retained earnings account because of the stock split? Theoretically, what will be the market price of the common stock immediately after the stock split?
Answer #1
a] Par value of preferred stock = 200000/10000 = $             20.00
b] Dividend per share of preferred stock = 20*6% = $               1.20
c] Average issue price of common stock = (1000000+500000)/100000 = $             15.00
e] # shares outstanding after stock split = 100000*2 = 200,000
No amount will be transferred from the retained earnings.
Market price, immediately after stock split, will be 42/2 = $             21.00